OpenStudy (anonymous):

Lila earns $8.75/h at her job. How many hours must Lila work to earn $105? Complete the equation to model the problem. Use h to represent the unknown value. ?=8.75*?

OpenStudy (anonymous):

105 = 8.75 * h -- divide both sides by 8.75 105/8.75 = h 12 = h she would have to work 12 hrs to make $ 105

OpenStudy (anonymous):

\[?=8.75\times?\]

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

no prob:)